# Faster Motion Planning via Restarts

### v1

@@ Abstract
Randomized methods such as PRM and RRT are widely used in motion planning. However, in some cases, their running-time suffers from inherent instability, leading to “catastrophic” performance even for relatively simple instances. We apply stochastic restart techniques, some of them new, for speeding up Las Vegas algorithms, that provide dramatic speedups in practice (a factor of $3$ [or larger] in many cases).
Our experiments demonstrate that the new algorithms have faster runtimes, shorter paths, and greater gains from multi-threading (when compared with straightforward parallel implementation). We prove the optimality of the new variants. Our implementation is open source, available on github, and is easy to deploy and use.
@@ 1 Introduction
SBMP  [^1] Sampling based motion planning. algorithms solve the robot motion planning problem, finding a valid motion of the robot between two given configurations, by randomly sampling the robot’s configuration space (c-space) consisting of all sequences of parametric values corresponding to robot’s configurations in the given environment (workspace), and using the samples to create a roadmap (graph) in that space.
@@
The probabilistic nature of these algorithms causes instability in the running time, as “incorrect” early decisions might lead to much longer exploration till a solution is found, see Figure 1.1.
@@
Figure 1.1: An illustration of two possible RRT search trees for a 2D point robot motion planning problem with a new extension in red. In (a), after adding the extension the algorithm must sample the purple region and then the orange region in order to find the passage. The extension in (b) results in a single, larger, region which will lead to the same result. In (c) we can see a possible result of an RRT run in which the first extension was chosen.
While some motion planning instances are relatively straightforward, robots with many degrees of freedom (DOF), can result in complicated and unintuitive motion planning tasks, with an associated c-space potentially containing narrow passages. Determining whether a given c-space contains such a narrow passage is computationally intractable [Rei79], making the estimation of an algorithm’s runtime (i.e., expectation or distribution) quite challenging without a battery of experiments. Narrow passages in the configuration space can lead to heavy-tailed runtime distributions. Such instances suffer from frequent catastrophic runtimes – far exceeding their mean or median. Counterintuitively, the motion planner might get lucky on these instances and solve them almost instantly. Unfortunately, deciding if a given instance has these properties is hard.
This problem is not unique to robot motion planning, and tools for addressing it have been developed and successfully used in several fields. See Section 2 for a short survey of related work. Consider any randomized algorithm alg, with running time $X$ (in seconds), where $X$ is a random variable, that always returns the correct answer. Randomized algorithms that never fail but their running time is a random variable are known as Las-Vegas algorithms.
@@
An example. Consider an algorithm alg with its running time $X$, such that ${\mathbb{P}}\left[X=1\right]=0.01$ (i.e., alg terminates after one second with probability $1/100$. Otherwise, it runs forever. (In practice, “forever” corresponds to cases where the algorithm just takes longer than what is acceptable ,e.g. minutes instead of seconds.) The optimal strategy here is obvious – run alg for one second, if it terminates, viola, a successful run was found. Otherwise, terminate it, and start a fresh run, with the same threshold. Clearly, the running time of this strategy has geometric distribution $\mathrm{Geom}(1/100)$, with expected running time of a $100$ seconds. A dramatic improvement over the original algorithm, which most likely would have run on forever. Note that even 32 or 64 threads are not a practical solution in this case.
@@
Of course, in most cases, the distribution $\mathcal{D}$ of the running time of alg is unknown. The challenge is thus to perform efficient simulation, that is, a restarting strategy, of alg without knowing $\mathcal{D}$. This problem was studied by Luby et al. [LSZ93], who presented an elegant strategy that runs in (expected) $O(\mathcal{O}\log\mathcal{O})$ time, where $\mathcal{O}$ is the expected running time of the optimal strategy for the given algorithm. They also prove that this bound is worst-case tight.
Many SBMP algorithms, RRT included, are probabilistically complete under reasonable assumptions [KSL+19, KSL+23, OCK24], implying they are Las-Vegas algorithms. Our purpose is to strategically choose restart times for the algorithm alg such that the strategy has a small (expected) running time, and has a “lighter” tail (i.e., one can prove a concentration bound on the running time until a successful run of the algorithm).
@@ 1.1 Contribution
In this paper, we present several strategies for preventing catastrophic runtimes in motion planning problems. We describe the time to live (TTL) simulation framework that all of the strategies share, the well-known strategy by Luby et al. [LSZ93] (see Section 4.1), and several new strategies for which we provide theoretical optimality analysis (see Section 4.2). We consider several different simulation models, starting with a simple “stop and restart” model, and a more involved schemes where the strategy creates several copies of the original algorithm and run them in parallel, or dovetail their execution, by pausing and resuming their execution, and some schemes in between.
@@
Our open source code implementation [AH25a] applies the presented strategies to algorithms provided via an executable. We demonstrate the benefits of using the different strategies with an extensive set of experiments on motion planning benchmarks (see Section 5). The experiments show drastically faster mean and median runtimes on tasks where RRT has an “unstable” runtime distribution, i.e. a distribution with a non-negligible probability of succeeding fast, coupled with a mean and median runtime heavily influenced by much longer runtimes. We also include the results of an experiment conducted on a task where RRT has a “stable” runtime distribution, meaning low variance, and verify that there is no advantage from using these strategies under such conditions. Furthermore, we show that the strategies also greatly benefit from parallelism, and, in certain tasks, also find shorter paths than those found by plain RRT. In particular, in some cases, there is factor six increase in the success rate of completing the motion planning task (on a single thread), see Table 5.2 (a). The effects are even more dramatic, in some cases, where the standard parallel-OR implementation completely fails to complete the task, while the new strategies succeed in the majority of runs, see Table 5.3 (a). Even in cases where the task is easier, and the parallel-OR complete all runs, the strategies exhibit a four to eight speedup, see Table 5.1.
@@
Paper layout. The paper is organized as follows. In Section 2 we review related work on stochastic resetting and SBMP algorithms. In Section 3 we discuss important preliminaries and introduce necessary notations. In Section 4 we describe the different strategies. We first explain the counter search strategy and the intuition behind it in Section 4.1, and then introduce our new strategies in Section 4.2. In Section 5 we present the experiments conducted and their results. Finally, we conclude in Section 6 with a discussion of our results.
2 Related work
2.1 Stochastic resetting
@@
Alt et al. [AGM+96] were the first to show how TTL strategies can be used to transform heavy-tailed Las Vegas algorithms to randomized algorithms with a small probability of a catastrophe. Inspired by that work  [^2] Published earlier, despite the years in the two references., Luby et al. [LSZ93] (mentioned above) gave provably optimal strategies for speeding up Las Vegas algorithms. They showed that if the distribution is given, a fixed TTL is optimal, and described an optimal strategy using a sequence of TTLs for unknown distributions.
@@
Luby and Ertel [LE94] studied the parallel version of Luby et al. [LSZ93]. They pointed out that computing the optimal strategy in this case seems to be difficult, but proved that under certain assumptions a fixed threshold strategy is still optimal. Furthermore, they show that running the strategy of Luby et al. on each CPU/thread yields a theoretically optimal threshold.
@@
The simulation techniques of Luby et al. are used in solvers for satisfiability [GKSS08], and constraint programming [vB06]. They are also used in more general search problems, such as path planning in graphs [MAQ23]. More widely, such strategies are related to stochastic resetting [EMS20], first passage under restart [PR17], and diffusion with resetting [TPS+20]. See [KR24] for a more comprehensive review of this method.
@@
2.2 Parallel-OR SBMP
A widely used strategy for alleviating the possibly heavy-tailed runtime distributions of SBMP algorithms is the parallel-OR model. First used in SBMP by Challou et al. [CGK93], this paradigm simply runs multiple copies of an algorithm using multithreading, and terminates as soon as the first thread succeeds. Challou et al. have shown speed-up super-linear in the number of processors, and described the mathematical reasoning of the paradigm which was later named “or parallel effect” [DSC13]. Carpin and Pagello [CP02] introduced parallel-OR RRT, which was shown to have an advantage in different scenarios by Aguinaga et al. [ABM08] and Devaurs et al. [DSC13]. Otte and Correll [OC13] introduced C-FOREST, which builds upon parallel-OR RRT and achieved super-linear speed-up for shortest path problems. Sun et al. [SPA15] have used parallel-OR RRT in order to quickly replan under uncertainty and achieve shorter paths and better probability of success.
@@ 3 Preliminaries
This section starts with several basic definitions, and continue by stating and proving several important claims and defining the profile of a Las-Vegas algorithm, which are required in order to follow the intuition behind the optimal strategy of Luby et al., referred to as counter search for reasons that will become apparent later, described in Section 4.1, and the optimality poofs of the new strategies in Section 4.2.
@@
Let $\mathbb{R}^{+\infty}$ denote the set of all real non-negative numbers including +$\infty$+ ∞. Let alg denote the given randomized algorithm, and assume its running time is distributed according to an (unknown) distribution $\mathcal{D}$ over $\mathbb{R}^{+\infty}$. Given a threshold $\alpha$, and a random variable $X\sim\mathcal{D}$, the $\alpha$-median of $\mathcal{D}$, denoted by $\mathcalb{m}_{\alpha}(X)=\mathcalb{m}_{\alpha}(\mathcal{D})$, is the minimum (formally, infimum) of $t$ such that ${\mathbb{P}}\left[X<t\right]\leq\alpha\text{ and }{\mathbb{P}}\left[X>t\right]\leq 1-\alpha.$ The median of $\mathcal{D}$, denoted by $\mathcalb{m}(\mathcal{D})$, is the $1/2$-median of $\mathcal{D}$.
@@
In the TTL strategy, we are given a prespecified sequence (potentially implicit and infinite) of TTLs $\mathcal{T}=t_{1},t_{2},t_{3}\ldots$, and run alg using these TTLs one after the other. Once the running time of an execution of alg exceeds the provided TTL, this copy of alg is killed and a new copy is started using the next TTL in the sequence.
@@
Observation 3.3.
The optimal strategy is always a fixed-TTL simulation
This is due to the memorylessness of the simulation process – the failure of the first round (which is independent of later rounds), does not change the optimal strategy to be used in the remaining rounds [^3] This is where we use the full information assumption – we gain no new information from knowing that the first run had failed.. Thus, whatever the threshold being used in the second round, should also be used in the first round, and vice versa. Thus, all rounds should use the same threshold. (For a formal proof of this, see Luby et al. [LSZ93]).
@@ 3.2 Algorithms and their profile
Lemma 3.4 (Full information settings [LSZ93, AH25b]).
@@
where $\qquad\qquad\text{where }R(t)=\frac{t}{{\mathbb{P}}\left[X\leq t\right]}.$
@@
$\mathrm{profile}\left(\texttt{alg}\right)=(1/\mathcalb{p},\mathcalb{t}),\qquad\text{where}\qquad\mathcalb{p}={\mathbb{P}}\left[X\leq\mathcalb{t}\right].$
@@
We show below, that, up a constant, the optimal strategy for alg with profile $(1/\mathcalb{p},\mathcalb{t})$ is to run the TTL simulation with threshold $\mathcalb{t}$. In expectation, this requires running alg $1/\mathcalb{p}$ times, each for $\mathcalb{t}$ time. This allows us later on to more easily prove the optimality of new strategies. The proof is mathematically neat, but the reader might benefit from skipping it on a first read.
@@
So assume $\beta>\tfrac{7}{8}$. Let $\mathcalb{m}=\mathbb{M}\!\left[X\;\middle|\;X\leq\Delta\right]$ and $\xi={\mathbb{E}}\!\left[X\;\middle|\;X\leq\Delta\right]$. By Markov’s inequality, we have that $\mathcalb{m}\leq 2\xi$. Namely, $f(\Delta)\geq\mathcalb{m}/2$. Observe that
@@ 4 Strategies
A significant portion of this section is dedicated to the counter search strategy, which we show is a valuable tool for SBMP algorithms. We thus not only describe the strategy, but also provide both intuition and mathematical analysis of the method in order to allow interested readers to more easily harness it.
After introducing and discussing counter search, two new strategies are presented. We start with strategies using only the stop/start operations while simulating an algorithm, but, rather than having a known infinite sequence of TTLs like counter search, our strategies draw a random TTL at each iteration based on carefully devised distributions.
Next we abandon the stop/start model, and allow pause/resume operations that give rise to two more strategies, one simulating concurrently running an algorithm at different speeds, and one caching a subset of earlier simulations of the algorithm, and resuming them rather than starting from scratch every time.

### v2

# Faster Sampling-Based Motion Planning via Restarts

@@ Abstract
Randomized methods such as PRM and RRT are widely used in motion planning. However, in some cases, their running time suffers from inherent instability, leading to “catastrophic” performance even for relatively simple instances. We propose stochastic restart techniques, some of them new, for speeding up randomized algorithms, that provide dramatic speedups in practice, a factor of 333 (or more in many cases).
Our experiments demonstrate that the new algorithms have faster runtimes, shorter paths, and greater gains from multi-threading (when compared with a straightforward parallel implementation). We prove the optimality of the new variants. Our implementation is open source, available on GitHub [AH25], and is easy to deploy and use.
@@ 1 Introduction
Sampling-based motion planning (SBMP) is a family of randomized algorithms used to solve the robot motion planning problem [Kav+96, Lav98]. Their probabilistic nature enables these methods to solve many instances of the problem despite its intractability [Rei79], making them widely used in robotics. Many SBMP algorithms, including RRT and PRM, are probabilistically complete under reasonable assumptions [Kle+19, Kle+23, OCK24], implying they are Las Vegas algorithms, that is, randomized algorithms that never fail but whose running time is a random variable.
@@
The probabilistic nature of these algorithms causes instability in the running time, as “incorrect” early decisions might lead to much longer exploration until a solution is found, see Figure 1.1. This instability may manifest as motion planning problems with heavy-tailed runtime distributions, meaning, intuitively, that the expected runtime increases significantly due to these high values.
@@
Figure 1.1: An illustration of two possible RRT search trees for a 2D point robot motion planning problem with a new extension in red. In (a), after adding the extension, the algorithm must sample the purple region and then the orange region to find the passage. The extension in (b) results in a single larger region, which will lead to the same result. In (c), we can see a possible result of an RRT run in which the first extension was chosen.
This problem of Las Vegas with heavy-tailed runtime distributions is not unique to robot motion planning, and tools for addressing it have been developed and successfully used in several fields.
@@
An example. Consider an algorithm alg with its running time $X$, such that ${\mathbb{P}}\left[X=1\right]=0.01$ (i.e., alg terminates after one second with probability 1/1001/1001 / 100. Otherwise, it runs forever. (In practice, “forever” corresponds to cases where the algorithm just takes longer than what is acceptable, e.g., minutes instead of seconds.) The optimal strategy here is obvious – run alg for one second, if it terminates, voilá, a successful run was found. Otherwise, terminate it and start a fresh run, with the same threshold. The running time of this strategy has a geometric distribution $\mathrm{Geom}(1/100)$, with expected running time of 100100100 seconds. A dramatic improvement over the original algorithm, which most likely would have run forever.
@@
The goal is to strategically choose restart times for the algorithm alg such that the strategy has a small (expected) running time, and has a “lighter” tail (i.e., one can prove a concentration bound on the running time until a successful run of the algorithm). Of course, in most cases, the distribution $\mathcal{D}$ of the running time of alg is unknown. The challenge is thus to perform efficient simulation, that is, a restarting strategy of alg without knowing $\mathcal{D}$. This problem was studied by Luby et al. [LSZ93], who presented an elegant strategy that runs in (expected) $O(\mathcal{O}\log\mathcal{O})$ time, where $\mathcal{O}$ is the expected running time of the optimal strategy for the given algorithm. They also prove that this bound is worst-case tight.
@@ 1.1 Contribution
In this paper, we present several strategies for preventing catastrophic runtimes in motion planning problems. We describe the time to live (TTL) simulation framework that all of the strategies share, the well-known strategy by Luby et al. [LSZ93] (see Section 3.1), and several new strategies for which we provide theoretical optimality analysis (see Section 3.2). We consider several different simulation models, starting with a simple “stop and restart” model, and more involved schemes where the strategy creates several copies of the original algorithm and runs them in parallel, or dovetails their execution, by pausing and resuming their execution, and some schemes in between.
@@
We demonstrate the benefits of using the different strategies with an extensive set of experiments on motion planning benchmarks (see Section 4). The experiments show drastically faster mean and median runtimes on tasks where RRT has an “unstable” runtime distribution, i.e., a distribution with a non-negligible probability of succeeding fast, coupled with a mean and median runtime heavily influenced by much longer runtimes. We also include the results of an experiment conducted on a task where RRT has a “stable” runtime distribution, meaning low variance, and verify that there is no advantage from using these strategies under such conditions.
Furthermore, we show that the strategies also greatly benefit from parallelism, and, in certain tasks, also find shorter paths than those found by plain RRT. In particular, in some cases, there is a factor six increase in the success rate of completing the motion planning task (on a single thread), see Table 4.2 (a). The effects are even more dramatic, in some cases, where the standard parallel-OR implementation, running multiple independent copies simultaneously, completely fails to complete the task, while the new strategies succeed in the majority of runs, see Table 4.3 (a). Even in cases where the task is easier, and the parallel-OR completes all runs, the strategies exhibit a four to eight speedup, see Table 4.1.
Our open source code implementation [AH25] applies the presented strategies to algorithms provided via an executable.
@@
Outline. The paper is organized as follows. In Section 1.2 we review related work on stochastic resetting and SBMP algorithms. In Section 2 we discuss important preliminaries and introduce necessary notations. In Section 3 we describe the different strategies. We first explain the counter search strategy and the intuition behind it in Section 3.1, and then introduce our new strategies in Section 3.2. In Section 4 we present the experiments conducted and their results. In Section 5 we provide guidance for applying restarting strategies in SBMP algorithms. Finally, we conclude in Section 6 with a discussion of our results.
1.2 Related work
1.2.1 Stochastic resetting
@@
Alt et al. [Alt+96] were the first to show how TTL strategies can be used to transform heavy-tailed Las Vegas algorithms to randomized algorithms with a small probability of a catastrophe. Inspired by that work, which was published earlier, despite the misleading publication dates, Luby et al. [LSZ93] (mentioned above) gave provably optimal strategies for speeding up Las Vegas algorithms. They showed that if the distribution is given, a fixed TTL is optimal, and described an optimal strategy using a sequence of TTLs for unknown distributions.
@@
Luby and Ertel [LE94] studied the parallel version of Luby et al. [LSZ93]. They pointed out that computing the optimal strategy in this case seems to be difficult, but proved that under certain assumptions, a fixed threshold strategy is still optimal. Furthermore, they show that running the strategy of Luby et al. on each CPU/thread yields a theoretically optimal threshold.
@@
The simulation techniques of Luby et al. are used in solvers for satisfiability [Gom+08], and constraint programming [Bee06]. They are also used in more general search problems, such as path planning in graphs [MAQ23]. More widely, such strategies are related to stochastic resetting [EMS20], first passage under restart [PR17], and diffusion with resetting [Tal+20]. See [KR24] for a more comprehensive review of this method.
@@
1.2.2 Parallel-OR SBMP
A widely used strategy for alleviating the possibly heavy-tailed runtime distributions of SBMP algorithms is the parallel-OR model. First used in SBMP by Challou et al. [CGK93], this paradigm simply runs multiple copies of an algorithm using multithreading, and terminates as soon as the first thread succeeds. Challou et al. have shown speed-up super-linear in the number of processors, and described the mathematical reasoning of the paradigm, which was later named “or parallel effect” [DSC13]. Carpin and Pagello [CP02] introduced parallel-OR RRT, which was shown to have an advantage in different scenarios by Aguinaga et al. [ABM08] and Devaurs et al. [DSC13]. Otte and Correll [OC13] introduced C-FOREST, which builds upon parallel-OR RRT and achieved super-linear speed-up for shortest path problems. Sun et al. [SPA15] have used parallel-OR RRT to quickly replan under uncertainty and achieve shorter paths and better probability of success.
[…]
@@ 2 Preliminaries
This section starts with several basic definitions, and continues by stating and proving several important claims and defining the profile of a Las Vegas algorithm. These are required to follow the intuition behind the optimal strategy of Luby et al., referred to as counter search for reasons that will become apparent later, described in Section 3.1, and the optimality proofs of the new strategies in Section 3.2.
@@
Let $\mathbb{R}^{+\infty}$ denote the set of all real non-negative numbers, including +$\infty$+ ∞. Let alg denote the given randomized algorithm, and assume its running time is distributed according to an (unknown) distribution $\mathcal{D}$ over $\mathbb{R}^{+\infty}$. Given a threshold $\alpha$, and a random variable $X\sim\mathcal{D}$, the $\alpha$-median of $\mathcal{D}$, denoted by $\mathcalb{m}_{\alpha}\left[X\right]=\mathcalb{m}_{\alpha}\left[\mathcal{D}\right]$, is the minimum (formally, infimum) of $t$ such that ${\mathbb{P}}\left[X<t\right]\leq\alpha\text{ and }{\mathbb{P}}\left[X>t\right]\leq 1-\alpha.$ The median of $\mathcal{D}$, denoted by $\mathcalb{m}\left[\mathcal{D}\right]$, is the 1/21/21 / 2-median of $\mathcal{D}$.
@@
In the TTL strategy, we are given a prespecified sequence (potentially implicit and infinite) of TTLs $\mathcal{T}=t_{1},t_{2},t_{3}\ldots$, and run alg using these TTLs one after the other. Once the running time of an execution of alg exceeds the provided TTL, this copy of alg is killed, and a new copy is started using the next TTL in the sequence.
@@
Observation 2.3.
The optimal strategy is always a fixed-TTL simulation.
This is due to the memorylessness of the simulation process – the failure of the first round (which is independent of later rounds), does not change the optimal strategy to be used in the remaining rounds [^1] This is where we use the full information assumption – we gain no new information from knowing that the first run had failed.. Thus, whatever the threshold being used in the second round should also be used in the first round, and vice versa. Thus, all rounds should use the same threshold. (For a formal proof of this, see Luby et al. [LSZ93]).
@@ 2.2 Algorithms and their profile
Lemma 2.4 (Full information settings [LSZ93]).
@@
$\qquad\qquad\text{where }R(t)=\frac{t}{{\mathbb{P}}\left[X\leq t\right]}.$
@@
$\mathrm{profile}\left(\texttt{alg}\right)=(1/\mathcalb{p},\mathcalb{t}),\qquad\text{where}\qquad\mathcalb{p}={\mathbb{P}}\left[X\leq\mathcalb{t}\right].$ (2.1)
@@
We show below that, up to a constant, the optimal strategy for alg with profile $(1/\mathcalb{p},\mathcalb{t})$ is to run the TTL simulation with threshold $\mathcalb{t}$. In expectation, this requires running alg $1/\mathcalb{p}$ times, each for $\mathcalb{t}$ time. This allows us later on to more easily prove the optimality of new strategies. The proof is mathematically neat, but the reader might benefit from skipping it on a first read.
@@
So assume $\beta>\tfrac{7}{8}$. Let $\mathcalb{m}=\mathcalb{m}\!\left[X\;\middle|\;X\leq\Delta\right]$ and $\xi={\mathbb{E}}\!\left[X\;\middle|\;X\leq\Delta\right]$. By Markov’s inequality, we have that $\mathcalb{m}\leq 2\xi$. Namely, $f(\Delta)\geq\mathcalb{m}/2$. Observe that
@@ 3 Strategies
A significant portion of this section is dedicated to the counter search strategy, which we show is a valuable tool for SBMP algorithms. We thus not only describe the strategy, but also provide both intuition and mathematical analysis of the method in order to allow interested readers to more easily harness it. This strategy is tried-and-tested in many applications, and our experimental results show its performance record extends to the motion planning problem.
After introducing and discussing counter search, two new strategies are presented. We start with strategies using only the stop/start operations while simulating an algorithm, but, rather than having a known infinite sequence of TTLs like counter search, our strategies draw a random TTL at each iteration based on carefully devised distributions. These strategies perform on par with the counter search strategy, but their random nature means they can be parallelized in a fully decentralized manner without creating a shared resource, making them easy to implement and use. However, the theoretical guarantees they provide are only with high probability.
Next, we abandon the stop/start model and allow pause/resume operations that give rise to two more strategies, one simulating concurrently running an algorithm at different speeds, and one caching a subset of earlier simulations of the algorithm and resuming them rather than starting from scratch every time. The idea behind this model is a very simple one – “waste not”. Runs of the algorithm whose TTL was reached without achieving success can be resumed to reach a higher TTL, thus reusing the computational effort invested. The experimental results reveal that these strategies are highly affected by hardware differences, and reusing the work is not enough to mitigate the overhead caused by the paused processes.